OpenStudy (anonymous):

Need help here please.... 7x+2y= 7/4 y=4x-1 The solution is ? (Type an ordered pair.)

OpenStudy (phi):

you could replace y in the first equation with 4x-1

OpenStudy (anonymous):

7x + 2y = 7/4 y = 4x - 1 substitute second eqn in the first eqn 7x + 2 (4x-1) = 7/4 solve for x

OpenStudy (anonymous):

Can you help me break it down, I am not great with fractions... (4x-1)= 3x?

OpenStudy (phi):

7x + 2 (4x-1) = 7/4 start with chan's equation: distribute the 2 over (4x-1) by multiplying everything inside the parens by 2

OpenStudy (anonymous):

8x-2=6x (that's what I got when I multiplied everything inside by 2)

OpenStudy (phi):

first: 2(4x-1) is 8x-2 which is what you got. next, you should get used to writing down the whole equation. so you would get 7x+2(4x-1)= 7/4 7x+8x-2 =7/4 now this is important: you can add 7x's to 8x's and get 15 x's but 8x-2 cannot be replaced with 6x (8x-2x is 6x) combine the x terms add +2 to both sides

OpenStudy (anonymous):

7x+2(4x-1)=7/4 7x+8x-2=7/4 15x-2=7/4

OpenStudy (anonymous):

am I on the right track?

OpenStudy (phi):

yes. now +2 on both sides

OpenStudy (anonymous):

15x-2=7/4 17x-4=7/4

OpenStudy (phi):

here is how: 15x-2=7/4 write +2 on both sides: 15x-2+2 = 7/4 +2 simplify the left side. and simplify the right side

OpenStudy (phi):

you can add or subtract numbers, so -2+2 becomes 0 on the right side, use a common denominator of 4: multiply 2 (2/1) by 4/4 then add the fractions

OpenStudy (anonymous):

15x-2+2=15x 7/4+2= 3 3/4

OpenStudy (anonymous):

I got the first one, but I think the second part of 7/4 +2 is wrong

OpenStudy (phi):

I would write it like this: 15x-2+2 = 7/4 +2 15x= 7/4 + 8/4 15x= 15/4 (add the tops after you have a common denominator) do you know what to do next to find x?

OpenStudy (anonymous):

7/4+8/4 =15/4 so, 15x=15/4 do I add 15 to both sides?

OpenStudy (phi):

you could but it would not be helpful what do you get if you multiplied both sides by 1/15: \[ \frac{1}{15}\cdot \frac{15x}{1} = \frac{1}{15}\cdot \frac{15}{4}\]

OpenStudy (anonymous):

15/60

OpenStudy (anonymous):

well, I did 1*15 then 15*4 to arrive at that answer

OpenStudy (phi):

yes. but you should write the entire equation

OpenStudy (anonymous):

1/15* 15/4=15/60

OpenStudy (phi):

the equation with x in it

OpenStudy (anonymous):

1/15*15x/1 that side I have no idea

OpenStudy (anonymous):

is it zero?

OpenStudy (phi):

that is the left side of the equation. you should write \[ \frac{1}{15}\cdot \frac{15x}{1} = \frac{15}{60} \] (this way you won't get lost as you take each step) But here is the "easy way": when you multiply fractions, you multiply top times top and bottom by bottom \[ \frac{1}{15}\cdot \frac{15x}{1} = \frac{1}{15}\cdot\frac{15}{4} \] you get \[ \frac{1\cdot 15 \cdot x}{15}= \frac{1\cdot15}{15 \cdot 4}\] order of multiplication can change (right?) 1*15= 15*1 that means this is the same as \[ \frac{15 \cdot x}{15}= \frac{15 \cdot 1}{15 \cdot 4}\] or "undoing" the fractions \[ \frac{15}{15} \cdot \frac{x}{1}= \frac{15 }{15}\cdot \frac{1}{4}\] what is 15/15 ?

OpenStudy (phi):

Once you learn this idea, you see that you can "cancel" \[ \frac{1}{\cancel{15}}\cdot \frac{\cancel{15}x}{1} = \frac{1}{\cancel{15}}\cdot \frac{\cancel{15}}{4} \]